OpenStudy (anonymous):

Need help with this one A) solve 4x^2=8 I think i got this one which is x=√2 B) find the x-intercepts of f(x) 4x^2-8 need help on understanding this one

OpenStudy (inkyvoyd):

brconnel, plus or minus sqrt(2) and, (sqrt(2),0) (-sqrt(2),0)

OpenStudy (anonymous):

would i use the quadratic equation for the intercepts though?

OpenStudy (chriss):

\[4x^{2}=8\] divide both sides by 4 \[x^{2}=4\] \[x=\sqrt{2},x=-\sqrt{2}\] To find the x-intercepts, you have to set y to 0. \[4x^{2}-8=0\] You can use the same results from part 8. \[x-intercepts\ are\ \left\{ (\sqrt{2},0),(-\sqrt{2},0) \right\}\]

OpenStudy (chriss):

part A, not part 8. sorry

OpenStudy (anonymous):

ok cool. Makes sense